A circular coil which has 100 turns has a radius o
Answer #1

The formula for flux at any time linked with the coil

\phi = N\pi r^2 B sin 2\pi ft

thus induced EMF can be derived as

E_{induced}=\frac{d\phi}{dt} = N2\pi^2 f r^2 B cos 2\pi ft

Hence maximum induced emf will be

E_{induced,max}=\frac{d\phi}{dt} = N2\pi^2 f r^2 B=454.33volt